# Face Forgery Detection with Elaborate Backbone

### v1

@@ 4.3.2 Adaptation Evaluation
Next, we evaluate the adaptation capability of various FFD methods by further fine-tuning FFD models using a few images from previously unseen common datasets, having initially trained these models on the FF++ dataset. Specifically, we conduct $10$-shot and $20$-shot experiments with an equal number of real and fake faces, fine-tuning the different FFD models for $2$ epochs with a learning rate of $52e-$ 5.
@@
4.4 Analysis of Backbone Fine-tuning Framework
TABLE V: Quantitative comparison of our framework with different components. Without UFM, we sum features from dual branches. When $\mathcal{L}_{EDU}$ is not used, $\mathcal{L}_{ce}$ is used instead. The results of our auxiliary branch (Ours_A) and both branches (Ours_M+A) are presented as reference.
@@ 4.4 Analysis of Backbone Fine-tuning Framework
Ours_A $79.83$ $68.29$ $68.97$ $72.36$ $67.96$
Ours_M+A $90.45$ $83.10$ $89.72$ $87.76$ $82.04$
Table V shows the quantitative comparison results of core components in our framework, as can be seen: (1) without $\mathcal{L}_{dec}$, the dual-branch backbone tends to learn similar forgery cues, resulting in only a slight improvement over BEiT v2’ under the fusion module with sum operation; (2) incorporating $\mathcal{L}_{dec}$ can better promote the generalization of the main branch, with the average AUC improving from $87.04$% to $88.14$%; (3) our framework with UFM provides dynamic weights for feature fusion based on the prediction uncertainties of the main and auxiliary branches, enhancing their competition and improving backbone performance; (4) $\mathcal{L}_{EDU}$ further enhances the model’s FFD capability by constraining prediction uncertainty and class probability.
@@
We further validate the flexibility of our competitive backbone fine-tuning framework by using backbones pre-trained with different learning approaches and datasets, including ViT with SL on ImageNet-1k (SL on 1K), MAE, MoCo v3, and BEiT v2. The quantitative comparison in Table VI shows that our framework with different backbones achieves around $2$% performance improvement, further demonstrating the robustness and potential of our competitive learning mechanism with the decorrelation constraint and uncertainty-based fusion module.
TABLE VI: Quantitative comparison of our framework with different backbones.
@@
Ours w/ ViT (SL on 1k) $97.10$ $83.41$ $78.33$ $75.09$ $78.94$
MAE’ (1k&Face) $98.37$ $85.85$ $78.04$ $78.07$ $80.65$
Ours w/ MAE’ (1k&Face) $97.97$ $85.91$ $79.05$ $81.16$ $82.04$
MoCo v3’ (1k&Face) $98.65$ $84.98$ $79.13$ $85.72$ $83.28$
Ours w/ MoCo v3’ (1k&Face) $98.48$ $88.02$ $79.17$ $89.19$ $85.46$
BEiT v2’ (1k&Face) $99.14$ $89.18$ $84.02$ $86.48$ $86.56$
Ours w/ BEiT v2’ (1k&Face) 99.36 90.46 84.90 90.97 88.78
TABLE VII: Quantitative comparison of pre-training on different scales of real faces.
Method Real Faces In-set Cross-set AUC $\uparrow$(%)
FF++ CDF DFDC FFIW Avg
BEiT v2’ 55w $99.14$ $89.18$ $84.02$ $86.48$ $86.56$
750w $99.21$ $90.76$ $86.77$ $88.82$ $88.78$
Ours 55w 99.36 $90.46$ $84.90$ 90.97 $88.78$
750w $99.13$ 90.99 86.79 $90.72$ 89.50
Moreover, to explain the performance improvement of our method, we visualize attention maps to show the model’s focus areas on the face and perform cluster analysis of the output in the feature space. For attention map visualization, we utilize Attention Rollout [32], which averages the attention weights across all heads and recursively multiplies them by the weight matrices of all layers in the backbone. Figure 6 illustrates that BEiT v2 better focuses on local facial components compared to the supervised learning approach. After real-face pre-training, BEiT v2’ effectively targets local forgery areas, while our model is even more precise, leading to more accurate and reliable detection results. For cluster analysis, we use t-SNE [125] to present the feature vectors of the last layer of the backbone. The isolated features of real and four forgery algorithms (Deepfakes, Face2Face, FaceSwap, NeuralTextures) in BEiT v2, as shown in Figure 7, indicate that BEiT v2 mainly learns method-specific types rather than general forgeries, making it difficult to generalize to a broader range of forgeries. In contrast, BEiT v2’ primarily distinguishes real and fake faces thanks to its robust facial representation capabilities from real face pre-training. Meanwhile, our model better clusters real faces and separates them from fake faces, indicating that it more effectively extracts implicit forgery cues for generalization.
Additionally, given the importance of specific knowledge from real faces for backbones in representing facial features, we introduce a larger-scale real-face dataset to validate its FFD performance. We build a $7$,$500$,$000$ real-face dataset based on MSCelebV1 [126], MegaFace [127], and LAION-Face [128] to further pre-train BEiT v2. The quantitative results in Table VII show that adding more real faces can further improve FFD performance, with our framework demonstrating the best results, despite the difficulty of further improvement at higher performance levels.
@@
Method Loss Threshold CDF to CDF&DFDC&FFIW DFDC to CDF&DFDC&FFIW FFIW to CDF&DFDC&FFIW Avg
CDF DFDC FFIW Avg CDF DFDC FFIW Avg CDF DFDC FFIW Avg
BEiT v2’ $\mathcal{L}_{ce}$ $0.5$ $77.61$ $72.00$ $76.00$ $75.20$ $77.61$ $72.00$ $76.00$ $75.20$ $77.61$ $72.00$ $76.00$ $75.20$ $75.20$
OT $83.40$ $75.40$ $78.00$ $78.93$ $80.69$ $76.02$ $76.60$ $77.77$ $81.47$ $75.00$ $80.20$ $78.89$ $78.53$
OT w/ U $83.01$ $75.70$ $79.60$ $79.44$ $81.85$ $75.80$ $80.20$ $79.28$ $82.05$ $75.76$ $80.40$ $79.40$ $79.37$
$\mathcal{L}_{EDU}$ $0.5$ $80.12$ $73.42$ $78.40$ $77.31$ $80.12$ $73.42$ $78.40$ $77.31$ $80.12$ $73.42$ $78.40$ $77.31$ $77.31$
OT $81.85$ $76.12$ $77.60$ $78.52$ $81.27$ $76.32$ $78.60$ $78.73$ $80.31$ $75.70$ $82.00$ $79.34$ $78.86$
OT w/ U $82.05$ $\underline{76.94}$ $\underline{82.00}$ $80.33$ $80.89$ $\underline{77.56}$ $\bm{84.00}$ $81.22$ $80.89$ $\underline{77.52}$ $\bm{85.60}$ $\underline{81.34}$ $\underline{80.96}$
Ours $\mathcal{L}_{ce}$ $0.5$ $83.40$ $75.70$ $81.40$ $80.17$ $83.40$ $75.70$ $81.40$ $80.17$ $\bm{83.40}$ $75.70$ $81.40$ $80.17$ $80.17$
OT $83.98$ $75.84$ $81.20$ $80.34$ $83.01$ $76.68$ $81.20$ $80.30$ $82.24$ $76.16$ $82.60$ $80.33$ $80.32$
OT w/ U $\underline{84.36}$ $76.54$ $80.00$ $80.30$ $\underline{83.59}$ $77.50$ $\underline{82.60}$ $\underline{81.23}$ $82.82$ $77.38$ $82.80$ $81.00$ $80.84$
OT $83.78$ $75.94$ $81.40$ $\underline{80.37}$ $81.66$ $77.06$ $82.20$ $80.31$ $82.63$ $76.76$ $84.20$ $81.20$ $80.63$
OT w/ U $\bm{85.33}$ $\bm{77.54}$ $\bm{83.80}$ $\bm{82.22}$ $\bm{84.75}$ $\bm{78.66}$ $\bm{84.00}$ $\bm{82.47}$ $\bm{83.40}$ $\bm{78.30}$ $\underline{84.80}$ $\bm{82.17}$ $\bm{82.29}$
4.5 Analysis of Threshold Optimization Mechanism
We conduct experiments to validate the critical role of the classification threshold $\tau$ in FFD final discriminant results. Our study includes the empirical value of $0.5$, using the optimal thresholds $\tau_{ot}$ from one cross-dataset to another, and our uncertainty-based threshold optimization mechanism (UTOM). Meanwhile, we compare $\mathcal{L}_{EDU}$ and $\mathcal{L}_{ce}$ to verify their impact on uncertainty and predicted probability.
[…]
TABLE IX: Quantitative comparison of different PAD methods from three source datasets to one target dataset.
Method Venue O&C&I to M O&M&I to C O&C&M to I I&C&M to O Avg
@@ 4.5 Analysis of Threshold Optimization Mechanism
DSCI [46] TIFS 2023 $\underline{5.48}$ $97.39$ $\underline{8.00}$ $\underline{97.50}$ $5.71$ $98.44$ $12.59$ $94.57$ $7.95$ $96.98$
UDG [47] ICCV 2023 $5.95$ $98.47$ $9.82$ $96.76$ $5.86$ $98.62$ $10.97$ $95.36$ $8.15$ $\underline{97.30}$
GAC-FAS [129] CVPR 2024 5.00 $97.56$ $8.20$ $95.16$ $\underline{4.29}$ $\underline{98.87}$ $\underline{8.60}$ $\underline{97.16}$ $\underline{6.52}$ $97.19$
Ours -- $5.95$ $\underline{98.52}$ 3.44 99.63 2.93 99.58 5.68 98.17 4.50 98.98
Figure 9: Visualization of attention maps from our PAD model on different target datasets. The model focuses on most facial regions in real faces, while highlighting inconsistent areas in fake faces.
@@ Acknowledgments
This work was supported by the National Natural Science Foundation of China (No. 62276249 and No. 62171421), the Strategic Priority Research Program of the Chinese Academy of Sciences (No. XDB0680000), the Fundamental Research Funds for the Central Universities (No. 202461014), the TaiShan Scholars Youth Expert Program of Shandong Province (No. tsqn202306096), the Beijing Nova Program (No. 20230484368), and the Youth Innovation Promotion Association CAS.

### v2

@@ 4.3.2 Adaptation Evaluation
Next, we evaluate the adaptation capability of various FFD methods by further fine-tuning FFD models using a few images from previously unseen common datasets, having initially trained these models on the FF++ dataset. Specifically, we conduct $10$-shot and $20$-shot experiments with an equal number of real and fake faces, fine-tuning the different FFD models for $2$ epochs with a learning rate of 2e52e-5.
@@
TABLE VI: Quantitative comparison of our framework with different components. Without UFM, we sum features from dual branches. When $\mathcal{L}_{EDU}$ is not used, $\mathcal{L}_{ce}$ is used instead. The results of our auxiliary branch (CFB_A) and both branches (CFB_M+A) are presented as reference.
@@
CFB_A $79.83$ $68.29$ $68.97$ $72.36$ $67.96$
CFB_M+A $90.45$ $83.10$ $89.72$ $87.76$ $82.04$
4.4 Analysis of Backbone Fine-tuning Framework
Table VI shows the quantitative comparison results of core components in our framework, as can be seen: (1) without $\mathcal{L}_{dec}$, the dual-branch backbone tends to learn similar forgery cues, resulting in only a slight improvement over FB under the fusion module with sum operation; (2) incorporating $\mathcal{L}_{dec}$ can better promote the generalization of the main branch, with the average AUC improving from 87.0487.04% to 88.1488.14%; (3) our framework with UFM provides dynamic weights for feature fusion based on the prediction uncertainties of the main and auxiliary branches, enhancing their competition and improving backbone performance; (4) $\mathcal{L}_{EDU}$ further enhances the model’s FFD capability by constraining prediction uncertainty and class probability.
@@ 4.4 Analysis of Backbone Fine-tuning Framework
We further validate the flexibility of our competitive backbone fine-tuning framework by using backbones pre-trained with different learning approaches and datasets, including ViT with SL on ImageNet-1k (SL on 1K), FaceMAE, FaceMoCo, and FaceBEiT (FB). The quantitative comparison in Table VII shows that our framework with different backbones achieves around 22% performance improvement, further demonstrating the robustness and potential of our competitive learning mechanism with the decorrelation constraint and uncertainty-based fusion module.
TABLE VII: Quantitative comparison of our framework with different backbones.
@@
Competitive ViT (SL on 1k) $97.10$ $83.41$ $78.33$ $75.09$ $78.94$
FaceMAE (1k&Face) $98.37$ $85.85$ $78.04$ $78.07$ $80.65$
Competitive FaceMAE (1k&Face) $97.97$ $85.91$ $79.05$ $81.16$ $82.04$
FaceMoCo (1k&Face) $98.65$ $84.98$ $79.13$ $85.72$ $83.28$
Competitive FaceMoCo (1k&Face) $98.48$ $88.02$ $79.17$ $89.19$ $85.46$
FaceBEiT (1k&Face) $99.14$ $89.18$ $84.02$ $86.48$ $86.56$
Competitive FaceBEiT (1k&Face) 99.36 90.46 84.90 90.97 88.78
@@
To explain the performance improvement, we visualize attention maps to highlight the model’s focus areas on the face and conduct cluster analysis in the feature space. For attention map visualization, we utilize Attention Rollout [32], which averages the attention weights across all heads and recursively multiplies them by the weight matrices of all layers in the backbone. Fig. 6 illustrates that BEiT v2 better focuses on local facial components compared to the supervised learning approach. After real-face pre-training, FB effectively targets local forgery areas, while our model is even more precise, leading to more accurate and reliable detection results. For cluster analysis, we use t-SNE [103] to present the feature vectors of the last layer of the backbone. The isolated features of real and four forgery algorithms (Deepfakes, Face2Face, FaceSwap, NeuralTextures) in BEiT v2, as shown in Fig. 7, indicate that BEiT v2’s limited ability to recognize general forgery cues constrains its generalization performance. In contrast, FB primarily distinguishes real and fake faces thanks to its robust facial representation capabilities from real face pre-training. Meanwhile, our model better clusters real faces and separates them from fake faces, indicating that it more effectively extracts implicit forgery cues for generalization.
4.5 Analysis of Facial Representation in FFD
Given the importance of specific knowledge from real faces for backbones in representing facial features, we introduce a larger-scale real-face dataset to validate its performance. We build a $7500k$ real-face dataset based on MSCelebV1 [39], MegaFace [78], and LAION-Face [138] to further pre-train FB, i.e., FB++. Table VIII shows that adding more real faces can further improve FFD performance, with our framework demonstrating the best results, despite the difficulty of further improvement at higher performance levels.
Previous studies [124, 30] propose that FFD models tend to focus on forgery-irrelevant features during training, such as facial identity and gender information, which hinders the FFD model’s generalization. Following CADDM [30], we freeze the model parameters and use its features to train a linear classifier for face recognition to analyze the retained facial identity features. The experimental results in Fig. 8 show that: (1) FB-pretrained (red dashed line) outperforms BEiT v2-pretrained (blue dashed line), demonstrating that pre-training on real faces improves facial representation and enhances performance in both face forgery detection and recognition tasks; (2) After backbone fine-tuning on the FFD dataset, BEiT v2-finetuned (blue solid line) increases its focus on facial identity features, while FB-finetuned (red solid line) significantly reduces this focus and improves FFD generalization, which suggests that the real-face pre-trained backbone captures comprehensive facial features rather than overfitting to facial identity information.
Figure 8: ID linear classification on frozen features of different backbone.
TABLE VIII: Quantitative comparison of pre-training on different scales of real faces.
Method Real Faces In-set Cross-set AUC $\uparrow$(%)
FF++ CDF DFDC FFIW Avg.
FB $550k$ $99.14$ $89.18$ $84.02$ $86.48$ $86.56$
FB++ $7500k$ $99.21$ $90.76$ $86.77$ $88.82$ $88.78$
CFB $550k$ 99.36 $90.46$ $84.90$ $90.97$ $88.78$
CFB++ $7500k$ $99.13$ $90.99$ $86.79$ $90.72$ $89.50$
CFB $w/$ SBI $550k$ $89.70$ $95.35$ $85.45$ $91.82$ $90.87$
CFB++ $w/$ SBI $7500k$ $91.96$ 95.46 88.66 92.31 92.14
@@ 4.5 Analysis of Facial Representation in FFD
Method Loss Threshold CDF to CDF&DFDC&FFIW DFDC to CDF&DFDC&FFIW FFIW to CDF&DFDC&FFIW Avg.
CDF DFDC FFIW Avg. CDF DFDC FFIW Avg. CDF DFDC FFIW Avg.
FB $\mathcal{L}_{ce}$ $\tau_{0.5}$ $77.61$ $72.00$ $76.00$ $75.20$ $77.61$ $72.00$ $76.00$ $75.20$ $77.61$ $72.00$ $76.00$ $75.20$ $75.20$
$\tau_{ot}$ $83.40$ $75.40$ $78.00$ $78.93$ $80.69$ $76.02$ $76.60$ $77.77$ $81.47$ $75.00$ $80.20$ $78.89$ $78.53$
$\tau_{ot}^{UTOM}$ $83.01$ $75.70$ $79.60$ $79.44$ $81.85$ $75.80$ $80.20$ $79.28$ $82.05$ $75.76$ $80.40$ $79.40$ $79.37$
$\mathcal{L}_{EDU}$ $\tau_{0.5}$ $80.12$ $73.42$ $78.40$ $77.31$ $80.12$ $73.42$ $78.40$ $77.31$ $80.12$ $73.42$ $78.40$ $77.31$ $77.31$
$\tau_{ot}$ $81.85$ $76.12$ $77.60$ $78.52$ $81.27$ $76.32$ $78.60$ $78.73$ $80.31$ $75.70$ $82.00$ $79.34$ $78.86$
$\tau_{ot}^{UTOM}$ $82.05$ $\underline{76.94}$ $\underline{82.00}$ $80.33$ $80.89$ $\underline{77.56}$ $\bm{84.00}$ $81.22$ $80.89$ $\underline{77.52}$ $\bm{85.60}$ $\underline{81.34}$ $\underline{80.96}$
CFB $\mathcal{L}_{ce}$ $\tau_{0.5}$ $83.40$ $75.70$ $81.40$ $80.17$ $83.40$ $75.70$ $81.40$ $80.17$ $\bm{83.40}$ $75.70$ $81.40$ $80.17$ $80.17$
$\tau_{ot}$ $83.98$ $75.84$ $81.20$ $80.34$ $83.01$ $76.68$ $81.20$ $80.30$ $82.24$ $76.16$ $82.60$ $80.33$ $80.32$
$\tau_{ot}^{UTOM}$ $\underline{84.36}$ $76.54$ $80.00$ $80.30$ $\underline{83.59}$ $77.50$ $\underline{82.60}$ $\underline{81.23}$ $82.82$ $77.38$ $82.80$ $81.00$ $80.84$
$\tau_{ot}$ $83.78$ $75.94$ $81.40$ $\underline{80.37}$ $81.66$ $77.06$ $82.20$ $80.31$ $82.63$ $76.76$ $84.20$ $81.20$ $80.63$
$\tau_{ot}^{UTOM}$ $\bm{85.33}$ $\bm{77.54}$ $\bm{83.80}$ $\bm{82.22}$ $\bm{84.75}$ $\bm{78.66}$ $\bm{84.00}$ $\bm{82.47}$ $\bm{83.40}$ $\bm{78.30}$ $\underline{84.80}$ $\bm{82.17}$ $\bm{82.29}$
4.6 Analysis of Threshold Optimization Mechanism
We validate the critical role of classification threshold $\tau$ in FFD final discriminant results. Our study includes empirical threshold $\tau_{0.5}$, optimal threshold $\tau_{ot}^{\text{UTOM}}$ computed on one cross-dataset and applied to other cross-datasets after adjusting prediction probabilities with our Uncertainty-based Threshold Optimization Mechanism (UTOM), and $\tau_{ot}$ obtained without probability adjustment as a comparison.
Table IX reports the quantitative comparisons, showing that $\tau_{ot}$ from one cross-dataset achieves better ACC on other cross-datasets compared to $\tau_{0.5}$, and $\tau_{ot}^{UTOM}$ performs the best. $\mathcal{L}_{EDU}$ also achieves better ACC compared to $\mathcal{L}_{ce}$, indicating that constraining uncertainty and predicted probability improves the model’s confidence in its predictions. Remarkably, $\tau_{ot}^{UTOM}$ with $\mathcal{L}_{EDU}$ achieves a significant ACC improvement over $\tau_{ot}$ with $\mathcal{L}_{ce}$ in both FB and CFB, with an average ACC increase of more than 22%, fully demonstrates the reliability of UTOM in FFD applications, as well as the importance of our fine-tuning framework with the uncertainty-based fusion module and $\mathcal{L}_{EDU}$.
Moreover, we gradually increase the number of visible cross-datasets (Source Dataset) to calculate $\tau_{ot}$ and $\tau_{ot}^{UTOM}$, and validate the applicability of our threshold optimization mechanism on fixed unseen cross-datasets (Target Dataset). Results in Table X show that the optimal threshold calculated across multiple datasets significantly improves FFD model accuracy, highlighting that leveraging more datasets with diverse forgeries helps compute better classification thresholds and enhances FFD model’s applicability. Besides, while UTOM proves effective overall, difficult samples with large $u$ may cause fake-to-real misclassifications. Severe class imbalance can compromise its robustness, requiring minority class re-sampling for cross-domain stability, which is further analyzed in Supp. File.
@@ 4.6 Analysis of Threshold Optimization Mechanism
Target Dataset Source Dataset Threshold
CDF WDF [141] DFD $\tau_{ot}$ $\tau_{ot}^{UTOM}$
DFDC+FFIW+DFDCP +AIFace [65] +UADFV [127] $76.52$ $77.00$
$76.52$ $77.68$
$\underline{78.28}$ 79.08
SadTalker SimSwap Inswap $\tau_{ot}$ $\tau_{ot}^{UTOM}$
FOMM+FaceDancer +MobileSwap+UniFace +LIA [111]+FSGAN [81] $89.94$ $92.08$
$90.00$ $\underline{92.74}$
$92.56$ 92.86
Figure 9: Visualization of attention maps from our PAD model on different target datasets. Highlighted areas indicate higher model attention.
TABLE XI: Quantitative comparison of different PAD methods from three source datasets to one target dataset.
Method Venue O&C&I to M O&M&I to C O&C&M to I I&C&M to O Avg.
@@
DSCI [73] TIFS 2023 $5.48$ $97.39$ $8.00$ $97.50$ $5.71$ $98.44$ $12.59$ $94.57$ $7.95$ $96.98$
UDG [69] ICCV 2023 $5.95$ $98.47$ $9.82$ $96.76$ $5.86$ $98.62$ $10.97$ $95.36$ $8.15$ $97.30$
GAC-FAS [59] CVPR 2024 $5.00$ $97.56$ $8.20$ $95.16$ $\underline{4.29}$ $\underline{98.87}$ $8.60$ $97.16$ $6.52$ $97.19$
GFPL-FAS [66] CVPR 2024 $\underline{3.09}$ $\underline{99.45}$ 2.56 $\underline{99.10}$ $5.43$ $98.41$ $\underline{3.33}$ $\underline{99.05}$ $\underline{3.60}$ $\underline{99.00}$
CFB-PAD - 2.62 99.74 $\underline{3.44}$ 99.63 2.93 99.58 3.30 99.53 3.07 99.62
@@ Acknowledgments
This work was supported by the National Natural Science Foundation of China (No. 62276249 and No. 62571503), the Strategic Priority Research Program of the Chinese Academy of Sciences (No. XDB0680000), the TaiShan Scholars Youth Expert Program of Shandong Province (No. tsqn202507108 and No. tsqn202306096), the Strategic Priority Research Program of CAS (No. XDB0680202), the Beijing Nova Program (No. 20230484368), and the Youth Innovation Promotion Association CAS.